# Dynamical breaking of inversion symmetry and strong second harmonic generation with nonlinear phonons

### v1

@@ Conclusion
In conclusion, we have described a new, symmetry breaking steady state for driven chiral, nonlinear phonons. This state is characterized by strong second harmonic generation and by the emergence of non-oscillating electric fields. These effects, being forbidden by the inversion symmetry of the underlying lattice, can serve as sharp experimental signatures of inversion symmetry breaking in the steady state and the effects presented in this manuscript. Beyond possible applications for second harmonic generation and rectification, the study of interactions of chiral phonons in the newly described state with electrons and other collective modes (e.g. magnons (kahana2024_chiral_rectification, )) offers an intriguing avenue for uncovering novel out-of-equilibrium correlated states and dynamical phase transitions.
@@ Acknowledgements.
We acknowledge comments by the anonymous referee of one of our previous publications (kiselev2019squid_spectroscopy, ), which, eventually, led to the research reported here, as well as useful discussions with Mark Rudner. We also thank Jonas F. Karcher who motivated us to work on this manuscript, by pointing out that “a Nature article about chiral phonons was on his newsfeed, and chiral phonons seem to be a hot topic” (juraschek2025chiral_phonons_rev, ). This project received funding from the Horizon Europe Marie Skłodowska-Curie Action program under Grant Agreement 101155351.

### v2

@@ Conclusion
In conclusion, we have described a new, symmetry breaking steady state for driven chiral, nonlinear phonons. This state is characterized by strong second harmonic generation and by the emergence of a ferroelectric response. These effects, being forbidden by the inversion symmetry of the underlying lattice, can serve as sharp experimental signatures of inversion symmetry breaking in the steady state and the effects presented in this manuscript. Beyond possible applications for second harmonic generation, rectification, and driven on-demand ferroelectricity, the study of interactions of chiral phonons in the newly described state with electrons and other collective modes (e.g. magnons (kahana2024_chiral_rectification, )) offers an intriguing avenue for uncovering novel out-of-equilibrium correlated states and dynamical phase transitions.
@@ Acknowledgements.
We acknowledge comments by the anonymous referee of one of our previous publications (kiselev2019squid_spectroscopy, ), which, eventually, led to the research reported here, as well as useful discussions with Mark Rudner. We also thank Jonas F. Karcher who motivated us to work on this manuscript, by pointing out that “a Nature article about chiral phonons was on his newsfeed (juraschek2025chiral_phonons_rev, ), and chiral phonons seem to be a hot topic”. This project received funding from the Horizon Europe Marie Skłodowska-Curie Action program under Grant Agreement 101155351.